Ammonia gas may be formed by the reaction of nitrogen gas and hydrogen gas. When 10.0g...

  1. Ammonia gas may be formed by the reaction of nitrogen gas and hydrogen gas. When 10.0g of nitrogen is combined with 5.0g of hydrogen:
  2. What is the limiting reagent?
  3. What is the theoretical yield?
  4. What is the mass of excess reagent remaining?
Answer #1

Ammonia is produced by reaction of nitrogen and and hydrogen.

N2 (g) + 3H2 (g) \rightarrow 2NH3 (g)

one mole of Nitrogen gas reacts with three moles Hydrogen gas.

for this reaction , N2 (g) :  3H2(g) = 1 : 3

We have, moles of Nitrogen gas (M.W. : 28 g ) = 10 g / 28 g/mol = 0.357 mole

  moles of Hydrogen gas (M.W. : 2 g ) = 5 g / 2 g/mol = 2.5 mole.

a. A limiting reagent is a substance in a chemical reaction, which is completely consumed during reaction, and its consumption stops reaction whatever amount of other reagents .

for this reaction required amounts    N2 (g) :  3H2(g) = 1 : 3

we have,   N2 (g) :  3H2(g) = 0.357 : 2.5 =  1 : 7

So, Nitrogen is limiting reagent.

b. Theoretical yield : It is yield obtained from complete conversion of limiting reagent .

We have 0.357 moles of limiting reagent Nitrogen , since 1 mol of N2 (g) reacts with 3 moles of H2(g) to produce 2 moles of Ammonia .

So, theoretical yield NH3 = 2*0.357 moles = 0.714 moles ( M.W. NH3 = 17 g /mol )

or  0.714 moles * 17 g /mol = 12.138 g.

c. Remaining Mass of excess reagent, H2(g) :

Moles of H2(g) reacted = 3*0.357 mole = 1.07 mole

Remaining = (2.5 - 1.07 ) moles = 1.43 moles

So, Remaining Mass of H2(g) = 2.86 g.